OpenStudy (anonymous):

Consider the curve given by y^2=4xy+1. Find dy/dx. I found (-2y)/(y-2x) = dy/dx. First, if you could validate that answer, then help me with this: Find all points on the curve where the slope of the tangent line is 2.

OpenStudy (anonymous):

I've started by setting \[-2y \div (y-2x)=2\], but I don't know where to go from there.

OpenStudy (bahrom7893):

use implicit differentiation: 2y(dy/dx)=4(x*(dy/dx)+y*1) + 0 2y(dy/dx) = 4x(dy/dx)+4y

OpenStudy (anonymous):

(2y)/(y-2x) = dy/dx. this is the answer

OpenStudy (bahrom7893):

2y(dy/dx)-4x(dy/dx)=4y Pull out the dy/dx: (dy/dx)(2y-4x)=4y dy/dx = 4y/(2y-4x) and through dividing everything by 2, you get waqassaddique's answer which is: dy/dx = 2y/(y-2x)

OpenStudy (anonymous):

Good, that's what I got. Can you help me with the second part? What are you supposed to do to get such an answer?

OpenStudy (bahrom7893):

Find all the points where M=2 means dy/dx=2

OpenStudy (anonymous):

Mhm.

OpenStudy (bahrom7893):

so: 2 = 2y/(y-2x) 1 = y/(y-2x) y-2x = y -2x = 0 x=0 Point: (0;f(0))

OpenStudy (bahrom7893):

f(0) = +/-1

OpenStudy (anonymous):

The y's just cancel out?

OpenStudy (anonymous):

\[y^2=4xy+1\] \[2yy'=4y+4xy'\] then algebra \[2yy'-4xy'=4y\] \[(2y-4x)y'=4y\] \[y'=\frac{4y}{2y-4x}=\frac{2y}{y-2x}\] is what i get

OpenStudy (anonymous):

yes and this is problem

OpenStudy (bahrom7893):

yea, just subtract y from both sides. Points are: (0;-1) and (0;1)

OpenStudy (anonymous):

yeah bahrom has the same

OpenStudy (anonymous):

Thank you all! :) I really appreciate it.

OpenStudy (bahrom7893):

yay, high five satellite!

OpenStudy (bahrom7893):

and u're welcome m.

OpenStudy (anonymous):

set = 2 and solve yes?

OpenStudy (anonymous):

Yes. bahrom just helped me with that. I needed help with the solving, but apparently the ys just cancel.